OpenStudy (twizttiez):

Which ordered pair is the solution to the system of equations? x+y=4 x-y=-6 A. (2, 2) B. (1, −7) C. (−1, 5) D. (−2, 4)

OpenStudy (anonymous):

D

OpenStudy (twizttiez):

How did you get D?

OpenStudy (anonymous):

nah my bad its C

OpenStudy (twizttiez):

How did you get that answer?

OpenStudy (anonymous):

-1+5=4 -1-5=-6

OpenStudy (twizttiez):

Ok thanks could you help on a few more?

OpenStudy (anonymous):

yes

OpenStudy (twizttiez):

Which ordered pair is the solution to the system of equations? 2x-5y=-6 x+2y=11 A. (3, 4) B. (5, 3) C. (7, 2) D. (15, 7)

OpenStudy (anonymous):

B

OpenStudy (twizttiez):

@TheCoque

OpenStudy (twizttiez):

I have about 3 more

OpenStudy (anonymous):

ok

OpenStudy (twizttiez):

Could you help?

OpenStudy (twizttiez):

Which ordered pair is the solution to the system of equations? -5x-3y=12 x-3y=-6 A. (3, −3) B. (−3, 1) C. (6, 4) D. (9, 5)

OpenStudy (twizttiez):

If you have ? marks popping up press F5

OpenStudy (anonymous):

B

OpenStudy (twizttiez):

100% i really need to pass

OpenStudy (twizttiez):

2 more lemme open a new ? so you can get more medals